# Spectral Representation-based Reinforcement Learning

### v1

@@ 8.2 Model-Based Reinforcement Learning
Model-based reinforcement learning (RL) learns a model of the environment dynamics from observed state transitions and rewards, and then leverages it to facilitate policy optimization (luo2024survey). One predominant approach treats the approximated model as a surrogate of the true environment, allowing the agent to generate synthetic trajectories through fictitious interaction (chua2018deep; luo2018algorithmic; janner2019trust; yu2020mopo). In essence, these synthetic trajectories augment the real-world experience by illustrating counterfactuals, depicting what might have happened if different actions were taken, thereby reducing the demand for trial-and-error in the real world. In terms of modeling, various generative models or architectures have been explored, spanning from Gaussian processes (kamthe2018data), neural network parameterized deterministic or stochastic models (chua2018deep; luo2018algorithmic; hansen2022temporal; hansen2023td; zhou2411dino), recurrent state space models (RSSMs) (hafner2019learning; hafner2019dream; hafner2020mastering; hafner2023mastering), transformers (chen2022transdreamer; micheli2022transformers), energy-based models (chen2024offline), and diffusion models (ding2024diffusion; alonso2024diffusion). Beyond the intricate design of the model architecture, the planning process itself also warrants careful design to mitigate the risk of generating unrealistic trajectories. This is because errors in state prediction can accumulate with long planning horizons, a phenomenon known as the compounding error. To manage this error, common strategies include branch rollout (janner2019trust), early truncation (kidambi2020morel), or reward penalty (yu2020mopo; sun2023model). In contrast, our proposed methods adopt a similar way to extract dynamics-informed representations from experiences but completely obviate the need for planning, thereby achieving greater computational efficiency.
@@
Apart from learning the dynamics model with task-specific interaction data, there has been a notable trend in adapting foundation models for dynamics prediction. This approach, first proposed as world models (ha2018recurrent), has been revitalized due to the advances in generative models. V-JEPA 2-AC (assran2025v) and DINO-world (baldassarre2025back) adopt similar approaches, by first pre-training visual encoders on large-scale video or image datasets, and subsequently post-training a predictive model with action-annotated datasets for control and planning. To even eliminate the need for such action-annotated datasets, Genie (bruce2024genie) infers latent actions from unlabelled videos and learns a world model modulated by such latent actions jointly. While substantial progress has been made on the visual pre-training, the dynamics modeling component remains comparatively underexplored. Spectral representations help fill this gap by offering a principled and diverse control-oriented perspective that can be paired with existing high-quality visual representations.
@@ 9 Closing Remarks
We believe our review and experiments reveal only a fraction of the potential held by spectral representations. Beyond online reinforcement learning, spectral representations provide a natural way to leverage offline (levine2020offline) or passive (ghosh2023reinforcement) interaction data, facilitating sample-efficient online adaptation. Unlike skill-based pre-training that often requires intricate algorithm design and is heavily dependent on the behavior policy, pre-training spectral representations only requires access to transition data, thus possessing superior transferability and generalization (ren2023spectral; wang2025learning). Furthermore, spectral representations can be readily extended to various RL scenarios, such as off-policy evaluation (hu2024primal), multi-agent RL (ren2024scalable), and goal-conditional RL (eysenbach2022contrastive), by identifying the spectral structure within the system dynamics and connecting them to the components in RL pipeline.
@@ Appendix A Online Exploration with Spectral Representations
Following guo2023provably, with the estimation of $\hat{\bm{\phi}}_{n}$ at the $n$-th episode, we introduce the bonus term $\hat{b}_{n}(s,a)$ to implement the principle of optimism in the face of uncertainty. However, since our representation lies in the kernel space $\mathcal{H}_{k}$, we follow yang2020function and define the bonus function in terms of the kernel:
@@ B.1 Implementation Details of Baseline Algorithms
[fujimoto2018addressing]. Due to its simplicity and strong empirical performance, we re-implemented TD3 in our code repository and built all spectral representation-based algorithms on TD3. The hyperparameters that we used to obtain the results are listed in Table 3. Note that we tuned certain design choices (such as target network update strategy) and hyperparameters based on our observation on the DMControl suite, and therefore, the final choice may slightly differ from the original implementation.
@@
[haarnoja2018soft]. Compared to TD3, SAC introduces the principle of maximum entropy into reinforcement learning and regularizes the traditional RL objective with entropy. Empirically, entropy regularization encourages the policy to explore the environment, thus potentially leading to improved performance. The hyperparameters of SAC are listed in Table 4. We enable the automatic tuning of the coefficient of the entropy term and set the target entropy to $\text{dim}(\mathcal{A})$ by default.
@@
[fujimoto2023sale]. TD7 incorporates several improvements over TD3, including representation learning by embedding prediction, 2) policy checkpointing, behavior cloning regularization for offline scenarios, and LAP replay buffer [fujimoto2020equivalence]. Different from how we leverage spectral representations, TD7 feeds the representations alongside with the raw observations and actions into the value networks. We re-implemented the TD7 algorithm for DMControl tasks and used the recommended hyperparameters in the author-provided implementation, which are listed in Table 5.
@@
[yarats2021mastering]. DrQ-V2 is a model-free reinforcement learning algorithm designed to efficiently solve complex control tasks by directly learning from raw pixel data. The core innovation of DrQ-V2 is to apply image augmentation to the raw observations of the agents to regularize the visual encoders, therefore preventing divergent optimization due the challenge of learning from pixels. Although the official implementation of DrQ-V2 uses different sets of hyperparameters for varying task difficulties, we used identical hyperparameters for all tasks in our experiments. The hyperparameters are listed in Table 6.
@@
[hafner2023mastering]. Dreamer-V3 is a general and highly performant model-based reinforcement learning algorithm. It core concept is to first learn an RSSM that predicts the future latent states based on past observations and actions, thus enabling efficient simulation within the latent space. The agent then learns optimal behaviors through imaginary interaction with this learned simulation, rather than through slow and costly real-world interaction. Such a model-based approach makes it incredibly data-efficient compared to model-free methods. For our experiments, we used existing PyTorch implementation of Dreamer-V3 provided in https://github.com/NM512/dreamerv3-torch and adapted it for our evaluation protocol.
@@
[hansen2023td]. TDMPC2 is another representative model-based reinforcement learning algorithm. Compared to Dreamer-V3, TDMPC2 learns a light-weight latent space dynamics model by predicting the representations of the next observations using the representation of current observations and actions. Another innovation of TDMPC2 is that, it incorporates model predictive control during sampling to actively explore the environments based on the current estimation of the model. This design significantly improves the undirectional exploration strategy used in common RL algorithms, and boosts sample efficiency. Similarly, we re-used the author provided PyTorch implementation and configurations in https://github.com/nicklashansen/tdmpc2, and only made modifications to adapt it for our evaluation protocol.
@@ B.2 Implementation Details of Spectral Representation-Based Algorithms
In our implementation, we extensively use ResidualMLP, a variant of MLP layers with residual connections, since we find it consistently demonstrates better plasticity and expressiveness, echoing the recent findings by nauman2024bigger and lee2024simba. A diagram of this architecture is provided in Figure 10.
@@
. LV-Rep uses a standard VAE architecture but with a learnable prior distribution. To prevent regularizing the posterior distribution with a bad prior in the initial stage of training, hafner2020mastering applies decoupled and asymmetric KL regularization for the posterior and prior. However, in our experiments, we found this provides marginal benefits to the performance, and therefore we a balanced regularization by default. The configurations of LV-Rep are listed in Table 8.

### v2

@@ 8.2 Model-Based Reinforcement Learning
Model-based reinforcement learning (RL) learns a model of the environment dynamics from observed state transitions and rewards, and then leverages it to facilitate policy optimization (Luo et al., 2024). One predominant approach treats the approximated model as a surrogate of the true environment, allowing the agent to generate synthetic trajectories through fictitious interaction (Chua et al., 2018; Luo et al., 2018; Janner et al., 2019; Yu et al., 2020). In essence, these synthetic trajectories augment the real-world experience by illustrating counterfactuals, depicting what might have happened if different actions were taken, thereby reducing the demand for trial-and-error in the real world. In terms of modeling, various generative models or architectures have been explored, spanning from Gaussian processes (Kamthe and Deisenroth, 2018), neural network parameterized deterministic or stochastic models (Chua et al., 2018; Luo et al., 2018; Hansen et al., 2022, 2023; Zhou et al., ), recurrent state space models (RSSMs) (Hafner et al., 2019b, a, 2020, 2023), transformers (Chen et al., 2022a; Micheli et al., 2022), energy-based models (Chen et al., 2024), and diffusion models (Ding et al., 2024; Alonso et al., 2024). Beyond the intricate design of the model architecture, the planning process itself also warrants careful design to mitigate the risk of generating unrealistic trajectories. This is because errors in state prediction can accumulate with long planning horizons, a phenomenon known as the compounding error. To manage this error, common strategies include branch rollout (Janner et al., 2019), early truncation (Kidambi et al., 2020), or reward penalty (Yu et al., 2020; Sun et al., 2023). In contrast, our proposed methods adopt a similar way to extract dynamics-informed representations from experiences but completely obviate the need for planning, thereby achieving greater computational efficiency.
@@
Apart from learning the dynamics model with task-specific interaction data, there has been a notable trend in adapting foundation models for dynamics prediction. This approach, first proposed as world models (Ha and Schmidhuber, 2018), has been revitalized due to the advances in generative models. V-JEPA 2-AC (Assran et al., 2025) and DINO-world (Baldassarre et al., 2025) adopt similar approaches, by first pre-training visual encoders on large-scale video or image datasets, and subsequently post-training a predictive model with action-annotated datasets for control and planning. To even eliminate the need for such action-annotated datasets, Genie (Bruce et al., 2024) infers latent actions from unlabelled videos and learns a world model modulated by such latent actions jointly. While substantial progress has been made on the visual pre-training, the dynamics modeling component remains comparatively underexplored. Spectral representations help fill this gap by offering a principled and diverse control-oriented perspective that can be paired with existing high-quality visual representations.
@@ 9 Closing Remarks
We believe our review and experiments reveal only a fraction of the potential held by spectral representations. Beyond online reinforcement learning, spectral representations provide a natural way to leverage offline (Levine et al., 2020) or passive (Ghosh et al., 2023) interaction data, facilitating sample-efficient online adaptation. Unlike skill-based pre-training that often requires intricate algorithm design and is heavily dependent on the behavior policy, pre-training spectral representations only requires access to transition data, thus possessing superior transferability and generalization (Ren et al., 2023b; Wang et al., 2025). Furthermore, spectral representations can be readily extended to various RL scenarios, such as off-policy evaluation (Hu et al., 2024), multi-agent RL (Ren et al., 2024), and goal-conditional RL (Eysenbach et al., 2022), by identifying the spectral structure within the system dynamics and connecting them to the components in RL pipeline.
@@ Appendix A Online Exploration with Spectral Representations
Following Guo et al. [2023], with the estimation of $\hat{\bm{\phi}}_{n}$ at the $n$-th episode, we introduce the bonus term $\hat{b}_{n}(s,a)$ to implement the principle of optimism in the face of uncertainty. However, since our representation lies in the kernel space $\mathcal{H}_{k}$, we follow Yang et al. [2020] and define the bonus function in terms of the kernel:
@@ B.1 Implementation Details of Baseline Algorithms
[Fujimoto et al., 2018]. Due to its simplicity and strong empirical performance, we re-implemented TD3 in our code repository and built all spectral representation-based algorithms on TD3. The hyperparameters that we used to obtain the results are listed in Table 3. Note that we tuned certain design choices (such as target network update strategy) and hyperparameters based on our observation on the DMControl suite, and therefore, the final choice may slightly differ from the original implementation.
@@
[Haarnoja et al., 2018]. Compared to TD3, SAC introduces the principle of maximum entropy into reinforcement learning and regularizes the traditional RL objective with entropy. Empirically, entropy regularization encourages the policy to explore the environment, thus potentially leading to improved performance. The hyperparameters of SAC are listed in Table 4. We enable the automatic tuning of the coefficient of the entropy term and set the target entropy to $\text{dim}(\mathcal{A})$ by default.
@@
[Fujimoto et al., 2023]. TD7 incorporates several improvements over TD3, including representation learning by embedding prediction, 2) policy checkpointing, behavior cloning regularization for offline scenarios, and LAP replay buffer [Fujimoto et al., 2020]. Different from how we leverage spectral representations, TD7 feeds the representations alongside with the raw observations and actions into the value networks. We re-implemented the TD7 algorithm for DMControl tasks and used the recommended hyperparameters in the author-provided implementation, which are listed in Table 5.
@@
[Yarats et al., 2021a]. DrQ-V2 is a model-free reinforcement learning algorithm designed to efficiently solve complex control tasks by directly learning from raw pixel data. The core innovation of DrQ-V2 is to apply image augmentation to the raw observations of the agents to regularize the visual encoders, therefore preventing divergent optimization due the challenge of learning from pixels. Although the official implementation of DrQ-V2 uses different sets of hyperparameters for varying task difficulties, we used identical hyperparameters for all tasks in our experiments. The hyperparameters are listed in Table 6.
@@
[Hafner et al., 2023]. Dreamer-V3 is a general and highly performant model-based reinforcement learning algorithm. It core concept is to first learn an RSSM that predicts the future latent states based on past observations and actions, thus enabling efficient simulation within the latent space. The agent then learns optimal behaviors through imaginary interaction with this learned simulation, rather than through slow and costly real-world interaction. Such a model-based approach makes it incredibly data-efficient compared to model-free methods. For our experiments, we used existing PyTorch implementation of Dreamer-V3 provided in https://github.com/NM512/dreamerv3-torch and adapted it for our evaluation protocol.
@@
[Hansen et al., 2023]. TDMPC2 is another representative model-based reinforcement learning algorithm. Compared to Dreamer-V3, TDMPC2 learns a light-weight latent space dynamics model by predicting the representations of the next observations using the representation of current observations and actions. Another innovation of TDMPC2 is that, it incorporates model predictive control during sampling to actively explore the environments based on the current estimation of the model. This design significantly improves the undirectional exploration strategy used in common RL algorithms, and boosts sample efficiency. Similarly, we re-used the author provided PyTorch implementation and configurations in https://github.com/nicklashansen/tdmpc2, and only made modifications to adapt it for our evaluation protocol.
@@ B.2 Implementation Details of Spectral Representation-Based Algorithms
In our implementation, we extensively use ResidualMLP, a variant of MLP layers with residual connections, since we find it consistently demonstrates better plasticity and expressiveness, echoing the recent findings by Nauman et al. [2024] and Lee et al. [2024]. A diagram of this architecture is provided in Figure 10.
@@
. LV-Rep uses a standard VAE architecture but with a learnable prior distribution. To prevent regularizing the posterior distribution with a bad prior in the initial stage of training, Hafner et al. [2020] applies decoupled and asymmetric KL regularization for the posterior and prior. However, in our experiments, we found this provides marginal benefits to the performance, and therefore we a balanced regularization by default. The configurations of LV-Rep are listed in Table 8.